OpenStudy (anonymous):

The 80th term of an arithmetic sequence is twice the 30th term. If the first term is 7, what is the 40th term?

OpenStudy (anonymous):

what do you know about arithmetic sequence so far?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

if you have treated arithmetic sequence, what are the formulas relating to an arithmetic sequence do you know?

OpenStudy (anonymous):

idk. we've learned about finding rules from a sequence before. does that make sense?

OpenStudy (anonymous):

do you know the following formula? \[\Large a_n =a_1+(n-1)d\] fro finding the nth term of an arithmetic sequence?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then you would have difficulties understanding my explanation...

OpenStudy (anonymous):

could you try to explain it to me?

OpenStudy (anonymous):

okay.. i would try to explain... since the 80th term is twice the 30th term, \[\Large a_{80}=2a_{30}\]

OpenStudy (anonymous):

so what does the d in the equation represent?

OpenStudy (anonymous):

i would break that down a bit using the formula i gave above... \[\Large a_1+(80-1)d=2(a_1+(30-1)d)\] you were told in the question that the first term of the sequence which is a(1) is 7... so \[\Large 7+79d=2(7+29d)\] \[\Large 7+29d=14+58d\] from there, what do you get when you solve for d??

OpenStudy (anonymous):

d represents the common difference between consecutive terms in the sequence.

OpenStudy (anonymous):

why did you subtract 1 from 30 and 80?

OpenStudy (anonymous):

that is what is in the formula..

OpenStudy (anonymous):

i posted the formula above..

OpenStudy (anonymous):

i got d=1/3

OpenStudy (anonymous):

i made a mistake! it is supposed to be \[\Large 7+79d=14+58d\]

OpenStudy (anonymous):

that's what i got and i solved for d.

OpenStudy (anonymous):

can you please post your working?

OpenStudy (anonymous):

7+(80-1)d=2(7+(30-1)d) 7+79d=2(7+29d) 7+79d=14+58d 21d=7 d=1/3

OpenStudy (anonymous):

yh..you are right! i made a mistake!

OpenStudy (anonymous):

okay now that you have found d from the information given in the question, you can now find the 40th term using the same formula i posted above.. \[\Large a_{40}=7+(40-1)\times \frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

a(40) simply represents the 40th term